OpenStudy (anonymous):

What is the radius of a communications satellite’s orbital path that is in a uniform circular orbit around Earth and has a period of exactly 24.0 hours

OpenStudy (badhi):

So they've given you the period of the circular motion of the satellite. From $$\omega=\frac{2\pi}T$$, you can find the angular velocity of the satellite. now, for a circular motion to happen,the centrifugal force(F) should be, $$F=mr\omega^2$$ Now this F should be generated from the gravitational field on the satellite. That gravitational force is, $$F=\frac{GmM}{r^2}$$ where M is the mass of the earth. equating these two forces, we get $$\begin {align*} mr\omega^2&=\frac{GmM}{r^2}\\ r\omega^2&=\frac{GM}{r^2}\\ r&=\sqrt[3]{\frac{GM}{\omega^2}}\\ r&=\sqrt[3]{\frac{GMT^2}{4\pi^2}} \end {align*}$$